OpenStudy (anonymous):

Solve 4x = 98. Round your answer to the nearest hundredth

ganeshie8 (ganeshie8):

\(\huge 4x = 98\) divide by 4 both sides \(\huge \frac{4x}{4} = \frac{98}{4}\) \(\huge \frac{\cancel4x}{\cancel4} = \frac{\cancel{98}^{49}}{\cancel{4}^2}\) \(\huge x = \frac{49}{2}\)

ganeshie8 (ganeshie8):

you need to long division to convert to decimal

ganeshie8 (ganeshie8):

u familar with that ?

OpenStudy (anonymous):

not rlly

ganeshie8 (ganeshie8):

okay. \(\huge \frac{49}{2}\) \(\huge \frac{48+1}{2}\) \(\huge \frac{48}{2} + \frac{1}{2}\)

ganeshie8 (ganeshie8):

hmm above steps make sense ?

ganeshie8 (ganeshie8):

48/2 = 24 1/2 = 0.5 you may learn long division @ http://www.mathsisfun.com/long_division3.html its a good page... explains very nicely. pls give a try